A ball is thrown upward from a height of 15 feet with an initial velocity of 5 ft/s. use the formula h(t) = -16t^2 + 5t + 15. How long will it take the ball to reach it's maximum height and how long will it take the ball to hit the ground?
Answer by Alan3354(69443)

A ball is thrown upward from a height of 15 feet with an initial velocity of 5 ft/s. use the formula h(t) = -16t^2 + 5t + 15. How long will it take the ball to reach it's [sic] maximum height and how long will it take the ball to hit the ground?
------------------
The max ht is the vertex of the parabola,
at t = -b/2a
t = -5/-32 = 5/32 seconds
-----------
It hits the ground then h(t) = 0
-16t^2 + 5t + 15 = 0
Solve for t. Ignore the negative solution.